# Causal machine learning for heterogeneous treatment effects in the presence of missing outcome data

### v1

@@ Appendix D Derivation of the EIF used within the mDR-learner and mEP-learner
$\leq$ $4\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{3}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{3}}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]}\|\pi-\pi_{0}\|^{2}\|\theta-\theta^{*}\|+$
$4\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{2}}\right)^{2}\right]}\|\pi-\pi_{0}\|\|\mu^{1}-\mu^{1}_{0}\|\|\theta-\theta^{*}\|+$
$4\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{2}}\right)^{2}\right]}\|\pi-\pi_{0}\|\|\mu^{0}-\mu^{0}_{0}\|\|\theta-\theta^{*}\|+$
$4\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}^{2}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{2}(1-\bar{\pi})^{2}}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]}\|\pi-\pi_{0}\|\|G-G_{0}\|\|\theta-\theta^{*}\|+$
$4\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}}\right)^{2}\right]}\|\mu^{1}-\mu^{1}_{0}\|\|G-G_{0}\|\|\theta-\theta^{*}\|+$
$4\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{3}\bar{\pi}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{3}(1-\bar{\pi})}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]}\|G-G_{0}\|^{2}\|\theta-\theta^{*}\|$
@@
$|D_{g}^{2}D_{\theta}$ $L(\theta^{*},\bar{g})[\theta-\theta^{*},g-g_{0},g-g_{0}]|\leq 4\left(\beta_{2,1}\|\pi-\pi_{0}\|^{2}\|\theta-\theta^{*}\|+\right.$
$\beta_{2,2}\|\pi-\pi_{0}\|\|\mu^{1}-\mu^{1}_{0}\|\|\theta-\theta^{*}\|+\beta_{2,3}\|\pi-\pi_{0}\|\|\mu^{0}-\mu^{0}_{0}\|\|\theta-\theta^{*}\|+$
$\beta_{2,4}\|\pi-\pi_{0}\|\|G-G_{0}\|\|\theta-\theta^{*}\|+\beta_{2,5}\|\mu^{1}-\mu^{1}_{0}\|\|G-G_{0}\|\|\theta-\theta^{*}\|+$
$\left.\beta_{2,6}\|G-G_{0}\|^{2}\|\theta-\theta^{*}\|\right)$
@@
$\beta_{2,1}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{3}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{3}}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]}$
$\beta_{2,2}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{2}}\right)^{2}\right]}$
$\beta_{2,3}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{2}}\right)^{2}\right]}$
$\beta_{2,4}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}^{2}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{2}(1-\bar{\pi})^{2}}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]}$
$\beta_{2,5}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}}\right)^{2}\right]}$
$\beta_{2,6}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{3}\bar{\pi}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{3}(1-\bar{\pi})}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]}$
@@
$|D_{g}^{2}D_{\theta}$ $L(\theta^{*},\bar{g})[\theta-\theta^{*},g-g_{0},g-g_{0}]|\leq 4\beta_{2}\|\theta-\theta^{*}\|\cdot\|g-g_{0}\|^{2}$
@@
$D_{\theta}^{2}L(\bar{\theta},\hat{g})[\hat{\theta}-\theta^{*},\hat{\theta}-\theta^{*}]\geq\lambda\|\hat{\theta}-\theta^{*}\|^{2}$
$\implies$ $\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|^{2}\leq L(\hat{\theta},\hat{g})-L(\theta^{*},\hat{g})-D_{\theta}L(\theta^{*},\hat{g})[\hat{\theta}-\theta^{*}]$
@@
$\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|^{2}\leq\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},\hat{g})[\hat{\theta}-\theta^{*}]$ (39)
@@
$\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|^{2}\leq\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]-\frac{1}{2}D_{g}^{2}D_{\theta}L(\theta^{*},\bar{g})[\hat{\theta}-\theta^{*},\hat{g}-g_{0},\hat{g}-g_{0}]$
@@
$-\frac{1}{2}D_{g}^{2}D_{\theta}$ $L(\theta^{*},\bar{g})[\hat{\theta}-\theta^{*},\hat{g}-g_{0},\hat{g}-g_{0}]\leq 2\left(\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\theta}-\theta^{*}\|+\right.$
$\beta_{2,2}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{1}-\mu^{1}_{0}\|\|\hat{\theta}-\theta^{*}\|+\beta_{2,3}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{0}-\mu^{0}_{0}\|\|\hat{\theta}-\theta^{*}\|+$
$\beta_{2,4}\|\hat{\pi}-\pi_{0}\|\|\hat{G}-G_{0}\|\|\hat{\theta}-\theta^{*}\|+\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|\|\hat{G}-G_{0}\|\|\hat{\theta}-\theta^{*}\|+$
$\left.\beta_{2,6}\|\hat{G}-G_{0}\|^{2}\|\hat{\theta}-\theta^{*}\|\right)$
@@
$\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|^{2}\leq$ $\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]+2\left(\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\theta}-\theta^{*}\|+\right.$
$\beta_{2,2}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{1}-\mu^{1}_{0}\|\|\hat{\theta}-\theta^{*}\|+\beta_{2,3}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{0}-\mu^{0}_{0}\|\|\hat{\theta}-\theta^{*}\|+$
$\beta_{2,4}\|\hat{\pi}-\pi_{0}\|\|\hat{G}-G_{0}\|\|\hat{\theta}-\theta^{*}\|+\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|\|\hat{G}-G_{0}\|\|\hat{\theta}-\theta^{*}\|+$
$\left.\beta_{2,6}\|\hat{G}-G_{0}\|^{2}\|\hat{\theta}-\theta^{*}\|\right)$
@@
$\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|^{2}\leq$ $\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]+\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{0}-\mu^{0}_{0}\|^{2}+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|^{4}+(\delta_{1}+\delta_{2}+\delta_{3}+\delta_{4}+\delta_{5}+\delta_{6})\|\hat{\theta}-\theta^{*}\|^{2}$

### v2

@@ Appendix D Derivation of the EIF used within the mDR-learner and mEP-learner
$\leq$ $4\left(\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{3}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{3}}\{Y-\bar{\mu}^{0}\}\right)^{2}(\pi-\pi_{0})^{4}\right]\right)^{1/2}\|\theta-\theta^{*}\|_{2}+$
$4\left(\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{2}}\right)^{2}(\pi-\pi_{0})^{2}(\mu^{1}-\mu^{1}_{0})^{2}\right]\right)^{1/2}\|\theta-\theta^{*}\|_{2}+$
$4\left(\operatorname{\mathbb{E}}\left[\left(\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{2}}\right)^{2}(\pi-\pi_{0})^{2}(\mu^{0}-\mu^{0}_{0})^{2}\right]\right)^{1/2}\|\theta-\theta^{*}\|_{2}+$
$4\left(\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}^{2}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{2}(1-\bar{\pi})^{2}}\{Y-\bar{\mu}^{0}\}\right)^{2}(\pi-\pi_{0})^{2}(G-G_{0})^{2}\right]\right)^{1/2}\|\theta-\theta^{*}\|_{2}+$
$4\left(\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}}\right)^{2}(\mu^{1}-\mu^{1}_{0})^{2}(G-G_{0})^{2}\right]\right)^{1/2}\|\theta-\theta^{*}\|_{2}+$
$4\left(\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{3}\bar{\pi}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{3}(1-\bar{\pi})}\{Y-\bar{\mu}^{0}\}\right)^{2}(G-G_{0})^{4}\right]\right)^{1/2}\|\theta-\theta^{*}\|_{2}$
$\leq$ $4\left(\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{3}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{3}}\{Y-\bar{\mu}^{0}\}\right)^{4}\right]^{1/2}\|\pi-\pi_{0}\|^{4}_{8}\right)^{1/2}\|\theta-\theta^{*}\|_{2}+$
$4\left(\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{2}}(\pi-\pi_{0})\right)^{4}\right]^{1/2}\|\mu^{1}-\mu^{1}_{0}\|_{4}^{2}\right)^{1/2}\|\theta-\theta^{*}\|_{2}+$
$4\left(\operatorname{\mathbb{E}}\left[\left(\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{2}}(\pi-\pi_{0})\right)^{4}\right]^{1/2}\|\mu^{0}-\mu^{0}_{0}\|_{4}^{2}\right)^{1/2}\|\theta-\theta^{*}\|_{2}+$
$4\left(\operatorname{\mathbb{E}}\left[\left(\left(\frac{AC}{\bar{G}^{2}\bar{\pi}^{2}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{2}(1-\bar{\pi})^{2}}\{Y-\bar{\mu}^{0}\}\right)(\pi-\pi_{0})\right)^{4}\right]^{1/2}\|G-G_{0}\|^{2}_{4}\right)^{1/2}\|\theta-\theta^{*}\|_{2}+$
$4\left(\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}}(\mu^{1}-\mu^{1}_{0})\right)^{4}\right]^{1/2}\|G-G_{0}\|^{2}_{4}\right)^{1/2}\|\theta-\theta^{*}\|_{2}+$
$4\left(\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{3}\bar{\pi}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{3}(1-\bar{\pi})}\{Y-\bar{\mu}^{0}\}\right)^{4}\right]^{1/2}\|G-G_{0}\|^{4}_{8}\right)^{1/2}\|\theta-\theta^{*}\|_{2}$
$=$ $4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{3}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{3}}\{Y-\bar{\mu}^{0}\}\right)^{4}\right]^{1/4}\|\pi-\pi_{0}\|^{2}_{8}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{2}}(\pi-\pi_{0})\right)^{4}\right]^{1/4}\|\mu^{1}-\mu^{1}_{0}\|_{4}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{2}}(\pi-\pi_{0})\right)^{4}\right]^{1/4}\|\mu^{0}-\mu^{0}_{0}\|_{4}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\left(\frac{AC}{\bar{G}^{2}\bar{\pi}^{2}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{2}(1-\bar{\pi})^{2}}\{Y-\bar{\mu}^{0}\}\right)(\pi-\pi_{0})\right)^{4}\right]^{1/4}\|G-G_{0}\|_{4}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}}(\mu^{1}-\mu^{1}_{0})\right)^{4}\right]^{1/4}\|G-G_{0}\|_{4}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{3}\bar{\pi}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{3}(1-\bar{\pi})}\{Y-\bar{\mu}^{0}\}\right)^{4}\right]^{1/4}\|G-G_{0}\|^{2}_{8}\|\theta-\theta^{*}\|_{2}$
$\leq$ $4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{3}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{3}}\{Y-\bar{\mu}^{0}\}\right)^{4}\right]^{1/4}\|\pi-\pi_{0}\|^{2}_{2}\|\theta-\theta^{*}\|+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{2}}(\pi-\pi_{0})\right)^{4}\right]^{1/4}\|\mu^{1}-\mu^{1}_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{2}}(\pi-\pi_{0})\right)^{4}\right]^{1/4}\|\mu^{0}-\mu^{0}_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\left(\frac{AC}{\bar{G}^{2}\bar{\pi}^{2}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{2}(1-\bar{\pi})^{2}}\{Y-\bar{\mu}^{0}\}\right)(\pi-\pi_{0})\right)^{4}\right]^{1/4}\|G-G_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}}(\mu^{1}-\mu^{1}_{0})\right)^{4}\right]^{1/4}\|G-G_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{3}\bar{\pi}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{3}(1-\bar{\pi})}\{Y-\bar{\mu}^{0}\}\right)^{4}\right]^{1/4}\|G-G_{0}\|^{2}_{2}\|\theta-\theta^{*}\|_{2}$
$\leq$ $4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{3}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{3}}\{Y-\bar{\mu}^{0}\}\right)^{4}\right]^{1/4}\|\pi-\pi_{0}\|^{2}_{2}\|\theta-\theta^{*}\|+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{2}}\right)^{8}\right]^{1/8}\|\pi-\pi_{0}\|_{2}\|\mu^{1}-\mu^{1}_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{2}}\right)^{8}\right]^{1/8}\|\pi-\pi_{0}\|_{2}\|\mu^{0}-\mu^{0}_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}^{2}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{2}(1-\bar{\pi})^{2}}\{Y-\bar{\mu}^{0}\}\right)^{8}\right]^{1/8}\|\pi-\pi_{0}\|_{2}\|G-G_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}}\right)^{8}\right]^{1/8}\|\mu^{1}-\mu^{1}_{0}\|_{2}\|G-G_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{3}\bar{\pi}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{3}(1-\bar{\pi})}\{Y-\bar{\mu}^{0}\}\right)^{4}\right]^{1/4}\|G-G_{0}\|^{2}_{2}\|\theta-\theta^{*}\|_{2}$
$\leq$ $4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{3}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{3}}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]^{1/2}\|\pi-\pi_{0}\|^{2}_{2}\|\theta-\theta^{*}\|+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{2}}\right)^{2}\right]^{1/2}\|\pi-\pi_{0}\|_{2}\|\mu^{1}-\mu^{1}_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{2}}\right)^{2}\right]^{1/2}\|\pi-\pi_{0}\|_{2}\|\mu^{0}-\mu^{0}_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}^{2}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{2}(1-\bar{\pi})^{2}}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]^{1/2}\|\pi-\pi_{0}\|_{2}\|G-G_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}}\right)^{2}\right]^{1/2}\|\mu^{1}-\mu^{1}_{0}\|_{2}\|G-G_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{3}\bar{\pi}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{3}(1-\bar{\pi})}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]^{1/2}\|G-G_{0}\|^{2}_{2}\|\theta-\theta^{*}\|_{2}$
@@
$|D_{g}^{2}D_{\theta}$ $L(\theta^{*},\bar{g})[\theta-\theta^{*},g-g_{0},g-g_{0}]|\leq 4\left(\beta_{2,1}\|\pi-\pi_{0}\|_{2}^{2}\|\theta-\theta^{*}\|_{2}+\right.$
$\beta_{2,2}\|\pi-\pi_{0}\|_{2}\|\mu^{1}-\mu^{1}_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+\beta_{2,3}\|\pi-\pi_{0}\|_{2}\|\mu^{0}-\mu^{0}_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$\beta_{2,4}\|\pi-\pi_{0}\|_{2}\|G-G_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+\beta_{2,5}\|\mu^{1}-\mu^{1}_{0}\|_{2}\|G-G_{0}\|_{2}\|\theta-\theta^{*}\|_{2}+$
$\left.\beta_{2,6}\|G-G_{0}\|_{2}^{2}\|\theta-\theta^{*}\|_{2}\right)$
@@
$\beta_{2,1}$ $=\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\hat{G}\hat{\pi}^{3}}\{Y-\hat{\mu}^{1}\}-\frac{(1-A)C}{\hat{G}(1-\hat{\pi})^{3}}\{Y-\hat{\mu}^{0}\}\right)^{2}\right]^{1/2}$
$\beta_{2,2}$ $=\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\hat{G}\hat{\pi}^{2}}\right)^{2}\right]^{1/2}$
$\beta_{2,3}$ $=\operatorname{\mathbb{E}}\left[\left(\frac{(1-A)C}{\hat{G}(1-\hat{\pi})^{2}}\right)^{2}\right]^{1/2}$
$\beta_{2,4}$ $=\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\hat{G}^{2}\hat{\pi}^{2}}\{Y-\hat{\mu}^{1}\}-\frac{(1-A)C}{\hat{G}^{2}(1-\hat{\pi})^{2}}\{Y-\hat{\mu}^{0}\}\right)^{2}\right]^{1/2}$
$\beta_{2,5}$ $=\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\hat{G}^{2}\hat{\pi}}\right)^{2}\right]^{1/2}$
$\beta_{2,6}$ $=\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\hat{G}^{3}\hat{\pi}}\{Y-\hat{\mu}^{1}\}-\frac{(1-A)C}{\hat{G}^{3}(1-\hat{\pi})}\{Y-\hat{\mu}^{0}\}\right)^{2}\right]^{1/2}$
@@
$|D_{g}^{2}D_{\theta}$ $L(\theta^{*},\bar{g})[\theta-\theta^{*},g-g_{0},g-g_{0}]|\leq 4\beta_{2}\|\theta-\theta^{*}\|_{2}\cdot\|g-g_{0}\|_{2}^{2}$
@@
$D_{\theta}^{2}L(\bar{\theta},\hat{g})[\hat{\theta}-\theta^{*},\hat{\theta}-\theta^{*}]\geq\lambda\|\hat{\theta}-\theta^{*}\|_{2}^{2}$
$\implies$ $\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|_{2}^{2}\leq L(\hat{\theta},\hat{g})-L(\theta^{*},\hat{g})-D_{\theta}L(\theta^{*},\hat{g})[\hat{\theta}-\theta^{*}]$
@@
$\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|_{2}^{2}\leq\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},\hat{g})[\hat{\theta}-\theta^{*}]$ (28)
@@
$\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|_{2}^{2}\leq\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]-\frac{1}{2}D_{g}^{2}D_{\theta}L(\theta^{*},\bar{g})[\hat{\theta}-\theta^{*},\hat{g}-g_{0},\hat{g}-g_{0}]$
@@
$-\frac{1}{2}D_{g}^{2}D_{\theta}$ $L(\theta^{*},\bar{g})[\hat{\theta}-\theta^{*},\hat{g}-g_{0},\hat{g}-g_{0}]\leq 2\left(\beta_{2,1}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{\theta}-\theta^{*}\|+\right.$
$\beta_{2,2}\|\hat{\pi}-\pi_{0}\|_{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}\|\hat{\theta}-\theta^{*}\|_{2}+\beta_{2,3}\|\hat{\pi}-\pi_{0}\|_{2}\|\hat{\mu}^{0}-\mu^{0}_{0}\|_{2}\|\hat{\theta}-\theta^{*}\|_{2}+$
$\beta_{2,4}\|\hat{\pi}-\pi_{0}\|_{2}\|\hat{G}-G_{0}\|_{2}\|\hat{\theta}-\theta^{*}\|_{2}+\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}\|\hat{G}-G_{0}\|_{2}\|\hat{\theta}-\theta^{*}\|_{2}+$
$\left.\beta_{2,6}\|\hat{G}-G_{0}\|_{2}^{2}\|\hat{\theta}-\theta^{*}\|_{2}\right)$
@@
$\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|_{2}^{2}\leq$ $\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]+2\left(\beta_{2,1}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{\theta}-\theta^{*}\|_{2}+\right.$
$\beta_{2,2}\|\hat{\pi}-\pi_{0}\|_{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}\|\hat{\theta}-\theta^{*}\|_{2}+\beta_{2,3}\|\hat{\pi}-\pi_{0}\|_{2}\|\hat{\mu}^{0}-\mu^{0}_{0}\|_{2}\|\hat{\theta}-\theta^{*}\|_{2}+$
$\beta_{2,4}\|\hat{\pi}-\pi_{0}\|_{2}\|\hat{G}-G_{0}\|_{2}\|\hat{\theta}-\theta^{*}\|_{2}+\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}\|\hat{G}-G_{0}\|_{2}\|\hat{\theta}-\theta^{*}\|_{2}+$
$\left.\beta_{2,6}\|\hat{G}-G_{0}\|_{2}^{2}\|\hat{\theta}-\theta^{*}\|_{2}\right)$
@@
$\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|_{2}^{2}\leq$ $\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]+\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|_{2}^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{\mu}^{0}-\mu^{0}_{0}\|_{2}^{2}+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|_{2}^{2}\|\hat{G}-G_{0}\|_{2}^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}^{2}\|\hat{G}-G_{0}\|_{2}^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|_{2}^{4}+(\delta_{1}+\delta_{2}+\delta_{3}+\delta_{4}+\delta_{5}+\delta_{6})\|\hat{\theta}-\theta^{*}\|_{2}^{2}$